OpenStudy (phagey):

You have 15 coins in your pocket that are either quarters (q) or nickels (n). They total $2.75. a. Write a system of linear equations that represents how many of each coin you have. b. How many of each coin do you have?

OpenStudy (phagey):

@zepdrix

OpenStudy (phagey):

someone help me please

ILovePuppiesLol (ilovepuppieslol):

@Jaynator495

OpenStudy (phagey):

help plz?

OpenStudy (zarkon):

what do you get when you add q and n?

ILovePuppiesLol (ilovepuppieslol):

qn?

OpenStudy (zarkon):

q=# quarters, n=#nickles

OpenStudy (zarkon):

q+n=?

ILovePuppiesLol (ilovepuppieslol):

idk 0___________0

OpenStudy (zarkon):

there are only nickels and quarters . So if you count up all the coins how many do you have?

OpenStudy (phagey):

8 quarters

OpenStudy (phagey):

and 15 nickels @Zarkon

OpenStudy (zarkon):

it says you only have 15 coins total "You have 15 coins in your pocket"

OpenStudy (phagey):

6 quarters and 9 nickels @Zarkon

OpenStudy (zarkon):

no...and aren't you supposed to set up a system of equations

OpenStudy (phagey):

@mathmale can you please help me

OpenStudy (mathmale):

"Zarkon q=# quarters, n=#nickles" this being algebra, you'll need to represent unknown quantities with letters. Logically enough, Zarkon has chosen q to represent the # of quarters and n to repr. the # of nickels. How many coins do you have in your pocket? What is their total worth?

OpenStudy (mathmale):

@phagey

OpenStudy (phagey):

i need only 15 coins and the total has to be 2.75 @mathmale

OpenStudy (phagey):

i cant figure out how many quarters or nickles though

OpenStudy (mathmale):

OK: Write equations!! The set of quarters and nickels adds up to $2.75 in value, and there are a total of 15 coins in the set. # of coins: q + can you finish this? Value of coins: $2.75 = q($0.25) + Can you finish this?

OpenStudy (mathmale):

You need two equations in two unknowns to solve this problem.

OpenStudy (phagey):

q+n=2.75

OpenStudy (phagey):

@mathmale

OpenStudy (phagey):

2.75=q(0.25)+n(0.5) @mathmale

OpenStudy (mathmale):

$2.75 pertains only to the value of the quarters and the value of the nickels. Your most recent equation is almost correct, but in regard to the value of a nickel you have a mistake. Could you find and fix that, please?

OpenStudy (mathmale):

"q+n=2.75" is incorrect. Why? Because q and n are counts (how many quarters, how many nickels, whereas $2.75 is VALUE / MONEY.

OpenStudy (phagey):

its not 0.5?

OpenStudy (mathmale):

No. That's 50 cents.

OpenStudy (mathmale):

What's the value of one nickel?

OpenStudy (phagey):

oh so 0.05? @mathmale

OpenStudy (mathmale):

Much better. Then, q + n = ? (0.25)q + (0.05)n = ?

OpenStudy (phagey):

so 2.75=q(0.25)+n(0.05) is the answer to the first part?

OpenStudy (mathmale):

It's one step on the way to solving this problem. By itself, that is not the solution, but it's a necessary part of the solution. What about the other equation?

OpenStudy (phagey):

what other equation the one about how many quarters and nickles?

OpenStudy (mathmale):

.25q + .05n = $2.75 is correct. The other equation necessary to solve this problem is: (yes, it's the one about how many quarters and how many nickels.)

OpenStudy (phagey):

thats the one part im having real trouble on @mathmale

OpenStudy (mathmale):

"You have 15 coins in your pocket that are either quarters (q) or nickels (n)." Total # of coins is 15. How many quarters have you? How many nickels? You must represent the number of each type of coin using a letter of the alphabet; use q and n. Write the equation that shows that the sum of the # of quarters and the # of nickels is 15.

OpenStudy (mathmale):

Hint: We've already done this. Just review the above dialogue and you'll find the equation in question. Again: One equation is for the value of the coins: .25q + 0.05n = $2.75 The other one is for the number of coins, ignoring their value. Where's that equation?

OpenStudy (phagey):

10 quarters and 5 nickles @mathmale

OpenStudy (mathmale):

Actually, we had this equation before. It's q + n = 15. The total number of coins is 15, whcih includes q quarters and n nickels. Thus you have the system of linear equations .25q + 0.05n = $2.75 q + n = 15 How would you solve tht for q and n? In your shoes I would multiply the first equation by 100, to eliminate the fractions .25 and 0.05.

OpenStudy (mathmale):

I'm sorry if this problem is hard for you. But for us to finish it, I need y our full attention.

OpenStudy (mathmale):

Remember, you can always ask questions to clear up things you don't understand.